# CuckooGraph: A Scalable and Space-Time Efficient Data Structure for Large-Scale Dynamic Graphs

### v1

@@ VI Conclusion
This paper proposes CuckooGraph, a storage scheme that is error-free, efficient, and can flexibly adapt to large-scale dynamic graphs. CuckooGraph consists of two key techniques: Transformation and Denylist. Transformation allows flexible spatial transformation based on the arrival of items to maximize memory space utilization; Denylist efficiently accommodates failed insertion items and cooperates with Transformation to speed up item processing. Theoretical analysis of CuckooGraph is provided. We evaluate CuckooGraph on multiple important graph tasks on four large-scale real-world datasets and one synthetic dense graph dataset. Experimental results show that CuckooGraph significantly outperforms the SOTA scheme.

### v2

@@ VI Conclusion
In this paper, we propose a novel data structure designed for large-scale dynamic graphs, called CuckooGraph, which includes two key techniques, Transformation and Denylist. Thanks to them, CuckooGraph can be flexibly resized based on actual operations to achieve memory efficiency while keeping few memory accesses to achieve fast processing speed without any prior knowledge of the upcoming graphs. Our mathematical analysis theoretically proves that CuckooGraph is time and space efficient. Our experimental results show that CuckooGraph significantly outperforms 4 state-of-the-art schemes. In particular, compared with Spruce, CuckooGraph achieves 32.66×32.$6$ faster insertion throughput while reducing memory space by about $32\%$, and 168.45×168.45$\times$168.45 × less running time on the SSSP task.